A single die is rolled five times, and the number of fours that come up is tallied. Find the probability of getting exactly two fours. (I want to know how to set it up to get me answer and the answer so that i can fill in the numbers that leads to that answer)
Answer by sudhanshu_kmr(1152)


Probability of getting 4 in single throw = 1/6
so, probability of net getting 4 in single throw = 5/6

probability of getting exactly two fours in five throw = 5C2* (1/6)^2 * (5/6)^3
( Binomial probability )
= 10 * 5^3 / (6^5)


I hope it will help you to understand the concept, feel free to ask if any doubt....